OpenStudy (anonymous):

How many solutions exist for the system y = 2x + 5 and y = -0.2x - 4? no solution one solution infinitely many solutions

hero (hero):

y = y 2x + 5 = -0.2x - 4 2x + .2x = -4 - 5 2.2x = -9 One solution

OpenStudy (anonymous):

y = 2x + 5 [1] y = -0.2x - 4 [2] [1]-[2] 0=2.02x+9 -9=2.02x x=-4.46 So y= 2*-4.46+5 =-3.92 So there is only one solution.

OpenStudy (anonymous):

:)